Andile earn 1250 per week and lindas earn 1500 per week calculate the ratio of lindas salary to andile salary in simplest form

1
Expert's answer
2022-01-25T11:21:45-0500

Andie earns 1250 per week, while Lindas earns 1500 per week

The ratio of Lindas to Andie is:


"\frac{1500}{1250}=\frac{30}{25}=\frac{6}{5}"

Thus, the ratio of Lindas to Andie to simplest form is


"6:5"